Use a quadratic Equation to find two numbers whose sum is 5 and whose product is -24
Answer by checkley75(3666)

X+Y=5 X=5-Y
XY=-24 NOW SUBSTITUTE (5-Y) FOR X AND SOLVE FOR Y
(5-Y)Y=-24
5Y-Y^2+24=0
Y^2-5Y-24=0
(Y-8)(Y+3)=0
Y-8=0
Y=8 ANSWER.
Y+3=0
Y=-3 ANSWER.
PROOF
-3+8=5
5=5
-3*8=-24
-24=-24